# TSPDiffuser: Diffusion Models as Learned Samplers for Traveling Salesperson Path Planning Problems

### v1

@@ 5.3 Results
Fig. 3a presents the relationships between SPL scores and computation times for all the methods. Despite being trained on synthetic data, TSPDiffuser effectively generalizes to real-world complex environments (a4 and a5), strongly indicating its transferability to real robots. It achieves high SPL scores with processing speeds that are almost 2x to 4x faster compared to the other baselines. Diffuser operates faster than TSPDiffuser, but its SPL score is significantly lower because a single sampling by diffusion models does not necessarily guarantee feasible or near-optimal paths. CPRM achieved comparable performance with much fewer nodes compared to PRM. Fig. 3b further elaborates on the computation times of TSPDiffuser and PRM. The construction of the roadmap is the main computational bottleneck. TSPDiffuser effectively tackles this problem by limiting the nodes and edges to be sampled around potential solution paths.
@@
We further investigated the scalability of TSPDiffuser with respect to the number of destinations: 20, 50, and 100, on the Standard environment. Based on the ablation study, we evaluated TSPDiffuser with $K=15$ to pursue more optimal solutions at the cost of required computation times. As summarized in Fig. 3d, TSPDiffuser ($K=15$) consistently outperforms the other baselines. It can find solutions with an SPL higher than $0.98$ in around 10 seconds for instances with 100 destinations, whereas PRM requires nearly 50 seconds to achieve the same level of performance.
@@
Fig. 4 visualizes typical solution paths and roadmaps obtained by each method. Compared to PRM, which uniformly samples nodes, data-driven planners such as CPRM, Diffuser, and TSPDiffuser concentrate more nodes around ground-truth paths. Diffuser sometimes failed to find feasible paths connecting all destinations without collisions. TSPDiffuser can find collision-free paths more reliably by generating multiple diverse paths to form a roadmap.
Figure 5: Failure case.
5.4 Limitations
Fig. 5 shows a simple case for a failure mode of TSPDiffuser. The diffusion model generates straight paths between two destinations despite the presence of obstacles in between. We speculate that this is partly due to the absence of highly skewed loop paths (as indicated by the black line) in the training data. Moreover, the current version of TSPDiffuser only supports the 2D state space, whereas practical robotics applications can sometimes involve more complex state spaces [14]. This requires extensions to the architectures for the instance encoder and noise estimation networks, which we seek in future work.
6 Conclusion

### v2

@@ V-C Results
Fig. 4a presents the relationships between SPL scores and computation times. Despite being trained on synthetic data, TSPDiffuser generalizes to real-world environments (a4 and a5) without additional training or parameter tuning, demonstrating its transferability to real-robot applications. It achieves high SPL scores with processing speeds that are almost 2x to 4x faster compared to the other baselines, even including model loading and running times on CPUs. This result shows that our approach is advantageous, especially in situations where the obstacle layout in the environment can change regularly. Diffuser operates faster than TSPDiffuser, but its SPL score is significantly lower because a single sampling by diffusion models does not necessarily guarantee feasible or near-optimal paths. CPRM achieved comparable performance with much fewer nodes compared to PRM. Fig. 4b elaborates on the computation times of TSPDiffuser and PRM. The construction of the roadmap is the main computational bottleneck. TSPDiffuser effectively tackles this problem by limiting the nodes and edges to be sampled around potential solution paths.
@@
We further investigated the scalability of TSPDiffuser with respect to the number of destinations: 20, 50, and 100, on the Standard environment. Based on the ablation study, we evaluated TSPDiffuser with $K=15$ to pursue more optimal solutions at the cost of required computation times. As summarized in Fig. 4d, TSPDiffuser ($K=15$) consistently outperforms the other baselines. It can find solutions with an SPL higher than $0.98$ in around 10 seconds for instances with 100 destinations, whereas PRM requires nearly 50 seconds to achieve the same level of performance. Note that as the number of destinations significantly increases, paths between them are more likely to be simple straight lines, where a more efficient edge validation method such as visibility roadmaps [41] would be effective.
@@
Fig. 5 visualizes typical solution paths and roadmaps obtained by each method. Compared to PRM, which uniformly samples nodes, data-driven planners such as CPRM, Diffuser, and TSPDiffuser concentrate more nodes around ground-truth paths. Diffuser sometimes failed to find feasible paths connecting all destinations without collisions. TSPDiffuser can find collision-free paths more reliably by generating diverse paths to form a roadmap.
V-D Limitations and Possible Extensions
TSPDiffuser currently supports only 2D state space. Extending our approach to higher dimensional state spaces is an interesting direction. If the environmental maps are still 2D, one can extend the denoising network to process higher-dimensional state trajectories, as already demonstrated by Diffuser [9]. If the environment is otherwise high-dimensional (e.g., 3D or more), the instance encoder would also need to be extended to accept higher-dimensional obstacle maps, similar to the approach used in Critical PRM [13]. Effective sampling around object boundaries (L9 in Alg. 1) would also become non-trivial for high-dimensional environments.
VI Conclusion